Solve for vo(t) using the Laplace methods.

6. Solve for vo(t) using Laplace methods. 4e-fu(t) A 1F (1 610) A 122 1 H - 312 103 0.00 2u(1) A
Answer #1

mesh mesh Gueurs TIF & buch) In laplace IH domain volts de In auch i curs ents. consider four mesh Iz = super me we can apply- S ) +224 + 4 + (5+1) -6=0 (5+3) I4 = 4 - 4 + 6 14 SC63 - SHC573) + 3 ] + sti 2 St3 74 = CBS 2-[5 Tu - 40 + 4674246 Iu = (4/